In closed, private economy the level of planned investment is 300 and the savings function is s=-100+0.25y.the equilibrium level of income and consumption respectively is
1
Expert's answer
2020-06-15T12:02:56-0400

Planned Investment = 300

Savings Function (S) S=-100 + 0.25Y

Therefore;

Under equilibrium condition Aggregate Demand = Aggregate Supply

AD = AS

Aggregate Income (Y) = Consumption (C) + Investment Expenditure (I) + Government Expenditure (G)

Y = C + I + G

Y = C + I + G + X - M

C = -100 + 0.25(Y - T)

Taxes (T) = 0

Government Spending(G) = 0

Investment Spending (I) = 300

Therefore;

Y = C = I + G + X - M

Y = -100 + 0.25(Y -0) + 300 + 0

Y = -100 + 0.25Y + 300

Y - 0.25Y = -100 + 300

0.75Y/0.75 = 200/0.75

Y = 266.70

Income = 266.70


Consumption (C)

C = -100 + 0.25(Y-T)

T = 0

Y = -100 + 0.25Y

Y - 0.25Y = -100

0.75Y/0.75 = -100/0.75

Y = -133.30

Consumption = 133.30